OpenStudy (anonymous):

HELP WILL MEDAL AND FAN Solve the following problem by working backward. Show and explain each step. Maria received a gift of cash from her grandmother. She put two-thirds of the money in her savings account, spent twenty-five dollars on a pair of jeans, and spent half the remaining money on a DVD. If Maria has eighteen dollars left over, how much money did she receive from her grandmother?

OpenStudy (anonymous):

What have you gotten so far?

OpenStudy (anonymous):

nothing i need help

OpenStudy (anonymous):

Well, let's go through this step by step. If she has 18 after spending half on the DVD, that means she had 2*18 before the DVD, right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

And 2*18 equals what?

OpenStudy (anonymous):

36

OpenStudy (anonymous):

Right. So now we take the amount spent on jeans and add it to 36.

OpenStudy (anonymous):

25+36=61

OpenStudy (anonymous):

Right. So since she spent put 2/3 in the savings account, that means we have 1/3 of the total right now at 61. So what will we multiply the 61 by to get the total?

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

Here, I'll write in equation form to make it a little easier. \[\frac{ 1 }{ 3 }*x=1\] What would x be?

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

idk if im right

OpenStudy (anonymous):

We're MULTIPLYING, not adding.

OpenStudy (anonymous):

oh then idk

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

Hint: It's the inverse of 1/3. Right, it would be 3.

OpenStudy (anonymous):

So 61*3=what?

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

$183

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (anonymous):

Yes, that would be it.